# DRAGON: Distributional Rewards Optimize Diffusion Generative Models

### v1

@@ 3.1 On-Policy Construction of ${\mathcal{D}}_{+}$ and ${\mathcal{D}}_{-}$
1:Query $f_{\theta}$ twice to get generations ${\mathcal{D}}_{1}=\{x_{11},\ldots,x_{1n}\}$ and ${\mathcal{D}}_{2}=\{x_{21},\ldots,x_{2n}\}$.11footnotemark: 1
$({\mathcal{D}}_{+}^{(0)},{\mathcal{D}}_{-}^{(0)})\leftarrow({\mathcal{D}}_{1},{\mathcal{D}}_{2})$ if $r_{\textrm{dist}}({\mathcal{D}}_{1})>r_{\textrm{dist}}({\mathcal{D}}_{2})$ else $({\mathcal{D}}_{2},{\mathcal{D}}_{1})$.
@@
4:     Swap the $i^{\text{th}}$ generation pair in ${\mathcal{D}}_{+}^{(i)}$ and ${\mathcal{D}}_{-}^{(i)}$ to form ${\mathcal{D}}_{+}^{{}^{\prime}(i)}$ and ${\mathcal{D}}_{-}^{{}^{\prime}(i)}$.
@@
For multi-GPU training parallelization, we broadcast all generations to each GPU and then only swap the indices originally present on each GPU. As a result, the initial sets $({\mathcal{D}}_{+}^{(0)},{\mathcal{D}}_{-}^{(0)})$ are identical across GPUs, but subsequent $({\mathcal{D}}_{+}^{(i)},{\mathcal{D}}_{-}^{(i)})$ may differ from different swapping indices. Even still, the copy of ${\mathcal{D}}_{+}$ per GPU is guaranteed to be as good as both ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$. The pseudocode in Appendix F includes this parallelization.
@@ 3.2 Learning From ${\mathcal{D}}_{+}$ And ${\mathcal{D}}_{-}$
We now optimize the generator parameters $\theta$ to make ${\mathcal{D}}_{\theta}$ attract ${\mathcal{D}}_{+}$ and repel ${\mathcal{D}}_{-}$, mathematically formulating this goal as minimizing the KL divergence $\mathrm{KL}({\mathcal{D}}_{+}\|{\mathcal{D}}_{\theta})$ and maximizing $\mathrm{KL}({\mathcal{D}}_{-}\|{\mathcal{D}}_{\theta})$. Intuitively, this means encouraging ${\mathcal{D}}_{\theta}$ to cover as much of ${\mathcal{D}}_{+}$ and as little of ${\mathcal{D}}_{-}$ as possible. Note that
[…]
@@
where $\pi_{\theta}(\cdot)$ denotes the likelihood for the model to generate a given example, and the integral is over the support of ${\mathcal{D}}_{+}$. Similarly, finding $\operatorname*{arg\,max}_{\theta}\mathrm{KL}({\mathcal{D}}_{-}\|{\mathcal{D}}_{\theta})$ is equivalent to finding $\operatorname*{arg\,min}_{\theta}\mathbb{E}_{x_{0}\sim{\mathcal{D}}_{+}}\left[\log\pi_{\theta}(x_{0})\right]$. Hence, our goal is equivalent to maximizing the log-likelihood for the model to generate examples in ${\mathcal{D}}_{+}$ and minimizing that of ${\mathcal{D}}_{-}$. For diffusion models, $\pi_{\theta}(\cdot)$ is implicit, making optimization more challenging compared to auto-regressive models. Despite this challenge, DRAGON can steer the likelihoods by leveraging specific loss functions, which include but are not limited to Diffusion-DPO loss (Wallace et al., 2024) and Diffusion-KTO loss (Li et al., 2024). Diffusion-DPO requires paired contrastive demonstrations, whereas Diffusion-KTO is more flexible and accepts unpaired ones. While Diffusion-DPO and Diffusion-KTO previously specialized in offline learning from large-scale human preference, they become components of an on-policy framework that optimizes arbitrary rewards when integrated into DRAGON. We provide mathematical details about these two loss functions in Section A.1 and empirically compare them in Section 5.2, where we also ablate between paired and unpaired demonstrations. In Section A.2, we discuss potential extensions beyond binary ${\mathcal{D}}_{+}$ and ${\mathcal{D}}_{-}$, compatibilities with alternative loss functions like GRPO (Shao et al., 2024) and DPOK (Fan et al., 2024), and relationships to reward-weighted regression.
@@
In total, DRAGON follows the illustration in Figure 1(a) and offers multiple advantages over existing reward optimization methods. First, DRAGON extends to distributional rewards like $r_{\textrm{dist}}$ that are hard/unstable to differentiate (e.g. FAD with large network backbones) via learning from contrastive demonstrations. Second, DRAGON allows for cross-modal supervision by learning from distributions rather than exact point-wise matches. This flexibility allows us to use cross-modal exemplar embeddings to construct rewards, even when the reference and generation modalities have substantial structural differences. In our experiments, we show that DRAGON can leverage text embeddings to improve music generation using only textual descriptions.
@@ 4.1 Instance-Wise Reward – Human Preference Dataset and Aesthetics Score Predictor
One of the most popular reward optimization tasks for content generation is aligning with human feedback, where human preference is the reward. To this end, human ratings of AI-generated examples provide relevant in-distribution guidance, and are thus more effective than ratings of human-created contents. Although open-source human preference datasets of AI image generations exist (Kirstain et al., 2023; Schuhmann, 2022; Schuhmann & contributors, 2022; Murray et al., 2012; Pressman et al., 2022), such resources are extremely rare for music. To demonstrate DRAGON’s ability to align music generations to human preferences, we collect a human rating dataset of AI-generated music and build a custom music aesthetics predictor model. This predictor serves dual purposes: a reward model for DRAGON to optimize, and an evaluation metric for evaluating DRAGON models trained for other reward functions.
Our human preference dataset, which we call Dynamo Music Aesthetics (DMA), consists of 800 prompts, 1,676 music pieces with various durations (total 15.97 hours), and 2,301 ratings from 63 raters on a scale of 1-5. The 1-5 rating scale makes our human feedback more fine-grained than binary pairwise comparison datasets such as (Kirstain et al., 2023). Details about this dataset and its collection are reported in Section B.1.
Our aesthetics predictor consists of a pre-trained CLAP audio encoder and a kernel regression prediction head, which we train on the DMA dataset. The textual prompt is not shown to the predictor. To determine model implementation details (e.g., music pre-processing, label normalization, CLAP embedding hop length) to optimize model performance on unseen data, we use a train/validation dataset split to perform an ablation study, which is presented in Section B.2. With model generalization verified, we remove the train/validation split and use all data to train the final predictor. A subjective test verified that generations with high predicted aesthetics scores indeed sound better than those with low scores, demonstrating more authentic instruments and better musicality. When the aesthetics score assigned by the predictor model is used as the reward function, DRAGON requires no additional music data.
@@ 4.2 Instance-to-Instance Reward – CLAP Score
We use CLAP score (Wu* et al., 2023), a popular music evaluation metric (Cideron et al., 2024; Bai et al., 2024; Hung et al., 2024), to demonstrate DRAGON’s capability to optimize instance-to-instance rewards. CLAP score is defined as the cosine similarity (clipped to be non-negative) between the CLAP embedding of a single generated audio instance and a single reference embedding. Leveraging CLAP’s unified cross-modal audio-text embedding space, we use the CLAP text embedding of the matching textual prompt as the reference for each audio generation. Intuitively, higher CLAP score means higher quality and semantic similarity. When maximizing CLAP score, DRAGON only requires a set of prompts and does not need any human-created music. When optimizing aesthetics score or CLAP score, both of which assign reward values to individual generations, ${\mathcal{D}}_{+}$ and ${\mathcal{D}}_{-}$ are constructed via pair-wise comparison. We find that DRAGON can improve overall generation quality by optimizing CLAP score.
@@ 4.3 Distribution-to-Distribution Reward – Full-Dataset FAD
We use the Fréchet audio distance (FAD) to demonstrate how we can accommodate reward signals that compare distributions or sets of generation outputs (audio) to corresponding target distributions (audio or text). FAD (lower is better) is one of the most commonly used metrics for evaluating music generation models (Kilgour et al., 2019). Intuitively, FAD represents the difference between a generated music distribution and a reference distribution (often human-created music) in an embedding space. Suppose that $\mu_{\theta}$ and $\mu_{\textrm{ref}}\in{\mathbb{R}}^{d}$ are respectively the means of the embeddings associated with generated and reference examples. Similarly, let $\Sigma_{\theta},\Sigma_{\textrm{ref}}\in{\mathbb{S}}_{+}^{d}$ denote the covariance matrices of the two distributions. FAD is computed as
$\textrm{FAD}\big{(}\mu_{\theta},\Sigma_{\theta},\mu_{\textrm{ref}},\Sigma_{\textrm{ref}}\big{)}\coloneqq\big{\|}\mu_{\theta}-\mu_{\textrm{ref}}\big{\|}_{2}^{2}+\textrm{Trace}\left(\Sigma_{\theta}+\Sigma_{\textrm{ref}}+\big{(}\Sigma_{\theta}^{\frac{1}{2}}\Sigma_{\textrm{ref}}\Sigma_{\theta}^{\frac{1}{2}}\big{)}^{\frac{1}{2}}\right).$ (3)
To minimize dataset FAD to match a reference distribution, we start by approximating the true generation distribution ${\mathcal{D}}_{\theta}$ with all generations in a training batch across all GPUs. Since full-dataset FAD assigns a reward value to a set of generations and does not rate each individual, ${\mathcal{D}}_{+}$ and ${\mathcal{D}}_{-}$ must be determined via Algorithm 1. I.e., the positive demonstrative set ${\mathcal{D}}_{+}$ is constructed with Algorithm 1 to have minimal dataset FAD. When multi-modal embedding spaces are used, the reference distribution can be in any supported modality. For example, when CLAP is used as the encoder, the reference can be either audio or text. When an audio embedding distribution is used as reference, DRAGON only requires the distribution’s mean and covariance. When the reference is a text embedding distribution, no audio data is needed for supervision.
@@ 4.4 Instance-to-Distribution Reward – Per-Song FAD
In addition to using FAD for distribution-to-distribution rewards, we also use FAD in an instance-to-distribution setting. While FAD is typically used to compare two distributions, it can also compare a single generation instance to a distribution by “bootstrapping” the instance. For music, we can split a generated waveform into shorter chunks and encode each chunk, forming a “per-song embedding distribution” (Gui et al., 2024). We can then use eq. 3 to compute the FAD between this single generation and the reference statistics. The reference statistics need not be per song and are computed using the entire reference dataset, and can again be in non-audio modalities if supported by the embedding space. Since per-song FAD is assigned to each example, DRAGON constructs the demonstration sets $({\mathcal{D}}_{+},{\mathcal{D}}_{-})$ via element-wise comparison (same as aesthetics and CLAP score optimization). In the literature, per-song FAD has been used to predict audio quality and identify dataset outliers (Gui et al., 2024). We show that with DRAGON, music generation models can improve generation quality by directly minimizing per-song FAD.
@@ 4.5 Reference-Free Distributional Reward – Embedding Diversity (Vendi Score)
The Vendi score, introduced in (Friedman & Dieng, 2023), is a diversity metric, for which a larger value means more diverse. Intuitively, a Vendi score of $v$ means that the diversity of a set of embeddings is similar to that of $v$ completely dissimilar vectors. To compute the Vendi score of given $n$ embeddings with dimension $d$ represented as a matrix $X\in{\mathbb{R}}^{n\times d}$, we first assemble an $n\times n$ positive semi-definite kernel matrix $K$. We use a linear kernel $K=\hat{X}\hat{X}^{\top}$, where $\hat{X}$ is obtained by normalizing $X$ so that each embedding has an $\ell_{2}$ norm of $1$. Next, we compute the eigenvalues of $K$, denoted as $\lambda_{1},\ldots,\lambda_{n}$. The Vendi score is then the eigenvalues’ exponentiated entropy:
$\textrm{Vendi}(X)\coloneqq\exp\big{(}-\sum_{i=1}^{n}\lambda_{i}\log\lambda_{i}\big{)}.$ (4)
During training, Vendi score is computed over generations in each training batch. We demonstrate directly improving Vendi with Algorithm 1, a result only possible because DRAGON operates on distributions.

### v2

@@ 3.1 On-Policy Construction of $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$
1:Query $f_{\theta}$ twice to get finite sets of generations ${\mathcal{D}}_{1}=\{x_{11},\ldots,x_{1n}\}$ and ${\mathcal{D}}_{2}=\{x_{21},\ldots,x_{2n}\}$.
[…]
@@
4:  At step $i$, swap the $i^{\text{th}}$ generation pair in ${\mathcal{D}}_{+}^{(i)}$ and ${\mathcal{D}}_{-}^{(i)}$ to form ${\mathcal{D}}_{+}^{{}^{\prime}(i)}$ and ${\mathcal{D}}_{-}^{{}^{\prime}(i)}$.
@@
For simplicity, we maintain equal sizes for ${\mathcal{D}}_{1}$, ${\mathcal{D}}_{2}$, $\mathcal{D}$$_{+}$, and $\mathcal{D}$$_{-}$ and define the stopping condition as a complete pass over ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$. Because ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$ are random roll-outs during training, we do not further randomize and swap the $i^{\text{th}}$ pair at the $i^{\text{th}}$ iteration. Each step of Algorithm˜1 is guaranteed to improve or maintain $\mathcal{D}$$_{+}$’s reward. When the loss function (to be discussed in Section˜3.2) requires paired demonstrations, we use the same conditioning but different random seeds to generate ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$. Subsequent swaps are also performed with same-conditioning pairs, ensuring paired elements in $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$ to provide direct contrast.
For multi-GPU training parallelization, we broadcast all generations to each GPU and then only swap the indices originally present on each GPU. As a result, the initial sets $({\mathcal{D}}_{+}^{(0)},{\mathcal{D}}_{-}^{(0)})$ are identical across GPUs, but subsequent $({\mathcal{D}}_{+}^{(i)},{\mathcal{D}}_{-}^{(i)})$ may differ from different swapping indices. Even still, the copy of $\mathcal{D}$$_{+}$ per GPU is guaranteed to be as good as both ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$. The pseudocode in Appendix˜F includes this parallelization.
@@ 3.2 Learning From $\mathcal{D}$$_{+}$ And $\mathcal{D}$$_{-}$
We now optimize the generator parameters $\theta$ to make ${\mathcal{D}}_{\theta}$ attract $\mathcal{D}$$_{+}$ and repel $\mathcal{D}$$_{-}$, mathematically formulating this conceptual goal as minimizing the KL divergence $\mathrm{KL}({\mathcal{D}}_{+}\|{\mathcal{D}}_{\theta})$ and maximizing $\mathrm{KL}({\mathcal{D}}_{-}\|{\mathcal{D}}_{\theta})$. Intuitively, this means encouraging ${\mathcal{D}}_{\theta}$ to cover as much of $\mathcal{D}$$_{+}$ and as little of $\mathcal{D}$$_{-}$ as possible. Note that
$\operatorname*{arg\,min}_{\theta}\mathrm{KL}({\mathcal{D}}_{+}\|{\mathcal{D}}_{\theta})\;=\;\operatorname*{arg\,max}_{\theta}\int\log\pi_{\theta}(x_{0})\ d{\mathcal{D}}_{+}(x_{0})\;=\;\operatorname*{arg\,max}_{\theta}\mathbb{E}_{x_{0}\sim{\mathcal{D}}_{+}}\left[\log\pi_{\theta}(x_{0})\right],$ (2)
@@
where $\pi_{\theta}(\cdot)$ denotes the likelihood for the model to generate a given example, and the integral is over the support of $\mathcal{D}$$_{+}$. Similarly, finding $\operatorname*{arg\,max}_{\theta}\mathrm{KL}({\mathcal{D}}_{-}\|{\mathcal{D}}_{\theta})$ is equivalent to finding $\operatorname*{arg\,min}_{\theta}\mathbb{E}_{x_{0}\sim{\mathcal{D}}_{+}}\left[\log\pi_{\theta}(x_{0})\right]$. Hence, our goal is equivalent to maximizing the log-likelihood for the model to generate examples in $\mathcal{D}$$_{+}$ and minimizing that of $\mathcal{D}$$_{-}$. Diffusion models pose a greater optimization challenge than auto-regressive ones because $\pi_{\theta}(\cdot)$ is implicit, making directly optimizing the KL-based conceptual objective impractical. To this end, DRAGON steers the likelihoods via surrogate loss functions, including but not limited to the Diffusion-DPO loss (Diffusion-DPO) and the Diffusion-KTO loss (Diffusion-KTO). Diffusion-DPO requires paired contrastive demonstrations, whereas Diffusion-KTO is more flexible and accepts unpaired ones. While Diffusion-DPO and Diffusion-KTO previously specialized in offline learning from large-scale preference datasets, they become components of an on-policy framework that optimizes arbitrary rewards when integrated into DRAGON, where $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$ continuously evolve with $\pi_{\theta}$. We provide mathematical details about these two loss functions in Section˜A.1 and empirically compare them in Section˜5.2, where we also ablate between paired and unpaired demonstrations. In Section˜A.2, we discuss potential extensions beyond binary $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$, compatibilities with alternative loss functions like GRPO (DeepSeekMath) and DPOK (DPOK), and relationships to reward-weighted regression.
@@
In total, DRAGON follows the illustration in Figure˜1(a) and offers multiple advantages over existing reward optimization methods. First, DRAGON extends to distributional rewards like $r_{\textrm{dist}}$ that are hard/unstable to differentiate (e.g. FAD with large network backbones) via learning from contrastive demonstrations. Second, DRAGON allows for cross-modal supervision by learning from distributions rather than exact point-wise matches. This flexibility allows us to use cross-modal exemplar embeddings to construct rewards, even when the reference and generation modalities have substantial structural differences. In our experiments, we show that DRAGON can leverage text embeddings to improve music generation using only textual descriptions.
@@ 4.1 Instance-Wise Reward – Human Preference Dataset and Aesthetics Score Predictor
One of the most popular reward optimization tasks for content generation is aligning with human feedback, where human preference is the reward. To this end, human ratings of AI-generated examples provide relevant in-distribution guidance, and are thus more effective than ratings of human-created contents. Although open-source human preference datasets of AI image generations exist (Pick-A-Pic; LAION-Aesthetics-Dataset; LAION-Aesthetics-Predictor-V2; AVA; SAC), such resources are extremely rare for music. To demonstrate DRAGON’s ability to align music generations to human preferences, we collect a human rating dataset of AI-generated music and build a custom music aesthetics predictor model. This predictor serves dual purposes: a reward model for DRAGON to optimize, and an evaluation metric for evaluating DRAGON models trained for other reward functions.
Our human preference dataset, which we call Dynamo Music Aesthetics (DMA), consists of 800 prompts, 1,676 music pieces with various durations (total 15.97 hours), and 2,301 ratings from 63 raters on a scale of 1–5. The 1–5 rating scale makes our human feedback more fine-grained than binary pairwise comparison datasets such as (Pick-A-Pic). DMA dataset and collection details are reported in Section˜B.1.
Our aesthetics predictor consists of a pre-trained CLAP audio encoder and a kernel regression prediction head, which we train on the DMA dataset. The textual prompt is not shown to the predictor. To determine model implementation details (e.g., music pre-processing, label normalization, CLAP embedding hop length) to optimize model performance on unseen data, we use a train/validation dataset split to perform an ablation study, which is presented in Section˜B.2. With model generalization verified, we remove the train/validation split and use all data to train the final predictor. A subjective test verified that generations with high predicted aesthetics scores indeed sound better than those with low scores, demonstrating more authentic instruments and better musicality. When the aesthetics score assigned by the predictor model is used as the reward function, DRAGON requires no additional music data.
@@ 4.2 Instance-to-Instance Reward – CLAP Score
We use CLAP score (laionclap2023), a popular music evaluation metric (MusicRL; ConsistencyTTA; TANGOFlux), to demonstrate DRAGON’s capability to optimize instance-to-instance rewards. CLAP score is defined as the cosine similarity (clipped to be non-negative) between the CLAP embedding of a single generated audio instance and a single reference embedding. Leveraging CLAP’s unified cross-modal audio-text embedding space, we use the CLAP text embedding of the matching textual prompt as the reference for each audio generation. Intuitively, higher CLAP score means higher quality and semantic similarity. When maximizing CLAP score, DRAGON only requires a set of prompts and does not need any human-created music. When optimizing aesthetics score or CLAP score, both of which assign reward values to individual generations, $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$ are constructed via pair-wise comparison. We find that DRAGON can improve overall generation quality by optimizing CLAP score.
@@ 4.3 Distribution-to-Distribution Reward – Full-Dataset FAD
We use the Fréchet audio distance (FAD) to demonstrate how we can accommodate reward signals that compare distributions or sets of generation outputs (audio) to corresponding target distributions (audio or text). FAD (lower is better) is one of the most commonly used metrics for evaluating music generation models (FAD). Intuitively, FAD represents the difference between a generated music distribution and a reference distribution (often human-created music) in an embedding space. Suppose that $\mu_{\theta}$ and $\mu_{\textrm{ref}}\in{\mathbb{R}}^{d}$ are respectively the means of the embeddings associated with generated and reference examples. Similarly, let $\Sigma_{\theta},\Sigma_{\textrm{ref}}\in{\mathbb{S}}_{+}^{d}$ denote the covariance matrices of the two distributions. FAD is computed as
$\textrm{FAD}\big(\mu_{\theta},\Sigma_{\theta},\mu_{\textrm{ref}},\Sigma_{\textrm{ref}}\big)\coloneqq\big\|\mu_{\theta}-\mu_{\textrm{ref}}\big\|_{2}^{2}+\textrm{Trace}\left(\Sigma_{\theta}+\Sigma_{\textrm{ref}}-2\big(\Sigma_{\theta}^{\frac{1}{2}}\Sigma_{\textrm{ref}}\Sigma_{\theta}^{\frac{1}{2}}\big)^{\frac{1}{2}}\right).$ (3)
To minimize dataset FAD to match a reference distribution, we start by approximating the true generation distribution ${\mathcal{D}}_{\theta}$ with all generations in a training batch across all GPUs. Since full-dataset FAD assigns a reward value to a set of generations and does not rate each individual, $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$ must be determined via Algorithm˜1. I.e., the positive demonstrative set $\mathcal{D}$$_{+}$ is constructed with Algorithm˜1 to have minimal dataset FAD. When multi-modal embedding spaces are used, the reference distribution can be in any supported modality. For example, when CLAP is used as the encoder, the reference can be either audio or text. When an audio embedding distribution is used as reference, DRAGON only requires the distribution’s mean and covariance. When the reference is a text embedding distribution, no audio data is needed for supervision.
@@ 4.4 Instance-to-Distribution Reward – Per-Song FAD
In addition to using FAD for distribution-to-distribution rewards, we also use FAD in an instance-to-distribution setting. While FAD typically compares two distributions, it can also compare a single generation instance to a distribution by “bootstrapping” the instance. For music, we can split a generated waveform into shorter chunks and encode each chunk, forming a “per-song embedding distribution” (FADTK). We can then use eq.˜3 to compute the FAD between this single generation and the reference statistics. The reference statistics need not be per song and are computed using the entire reference dataset, and can again be in non-audio modalities if supported by the embedding space. Since per-song FAD is assigned to each example, DRAGON constructs the demonstration sets $({\mathcal{D}}_{+},{\mathcal{D}}_{-})$ via element-wise comparison (same as aesthetics and CLAP score optimization). In the literature, per-song FAD has been used to predict audio quality and identify dataset outliers (FADTK). We show that with DRAGON, music generation models can improve generation quality by directly minimizing per-song FAD.
@@ 4.5 Reference-Free Distributional Reward – Embedding Diversity (Vendi Score)
The Vendi score (Vendi) is a diversity metric, for which a larger value means more diverse. Intuitively, a Vendi score of $v$ means that the diversity of a set of embeddings is similar to that of $v$ completely dissimilar vectors. To compute the Vendi score of given $n$ embeddings with dimension $d$ represented as a matrix $X\in{\mathbb{R}}^{n\times d}$, we first assemble an $n\times n$ positive semi-definite kernel matrix $K$. We use a linear kernel $K=\hat{X}\hat{X}^{\top}$, where $\hat{X}$ is obtained by normalizing $X$ so that each embedding has an $\ell_{2}$ norm of $1$. Next, we compute the eigenvalues of $K$, denoted as $\lambda_{1},\ldots,\lambda_{n}$. The Vendi score is then the eigenvalues’ exponentiated entropy:
$\textrm{Vendi}(X)\coloneqq\exp\big(\hskip-2.84526pt-\sum_{i=1}^{n}\lambda_{i}\log\lambda_{i}\big).$ (4)
During training, Vendi score is computed over generations in each training batch. We demonstrate directly improving Vendi with Algorithm˜1, a result only possible because DRAGON operates on distributions.